# Two-Stage Quranic QA via Ensemble Retrieval and Instruction-Tuned Answer Extraction

### v1

@@ I Introduction
The Holy Qur’an, revealed over 1,400 years ago, remains the primary source of guidance for over 1.8 billion Muslims worldwide. Beyond its religious significance, the Qur’an represents a masterpiece of Classical Arabic literature, containing profound linguistic, historical, and ethical insights that continue to be studied by scholars across multiple disciplines [4]. However, accessing and understanding specific Quranic content in response to natural language queries presents significant computational challenges that have only recently begun to be addressed by the natural language processing community [5].
Question Answering (QA) systems have emerged as a cornerstone technology for information access across diverse domains, from general knowledge retrieval to specialized applications in legal, medical, and educational contexts [1, 2]. The development of effective QA systems for religious texts, particularly the Qur’an, represents both a meaningful application with substantial social impact and a challenging technical problem that pushes the boundaries of current Arabic NLP capabilities [5].
Quranic Question Answering presents unique challenges that distinguish it from conventional QA tasks. First, the linguistic complexity of Classical Arabic—with its rich morphology, complex syntax, and archaic vocabulary—poses significant difficulties for models primarily trained on Modern Standard Arabic or contemporary dialects [7]. Second, the semantic richness and multi-layered meanings inherent in Quranic verses require sophisticated understanding that goes beyond surface-level text matching. Third, the limited availability of large-scale, high-quality annotated datasets for Quranic content constrains the development and evaluation of machine learning models [18, 24, 26]. Finally, the cultural and religious sensitivity of the domain demands high precision and reliability, as incorrect interpretations could have serious implications for users seeking religious guidance [3].
Recent advances in transformer-based language models have shown promising results for Arabic NLP tasks [9, 10]. However, these models face particular challenges when applied to Classical Arabic texts, where the linguistic gap between training data and target domain significantly impacts performance [11]. While ensemble methods have proven effective in improving robustness across various NLP tasks [29], their application to low-resource, high-complexity domains like Quranic QA remains underexplored [30]. Similarly, the emergence of instruction-tuned large language models offers new possibilities for few-shot learning in specialized domains, yet their effectiveness for Classical Arabic text understanding has not been thoroughly investigated [8, 12, 13].
@@
This paper addresses the identified limitations through a comprehensive two-stage framework that combines strategic dataset enhancement with complementary modeling approaches. Our key insight is that the challenges of passage retrieval and answer extraction in Quranic QA require different solutions: while retrieval benefits from ensemble methods applied to fine-tuned domain-specific models, extraction performs better with instruction-tuned large language models using carefully designed prompting strategies.
@@ II-C Ensemble Approaches
Ensemble methods have demonstrated significant benefits in QA tasks. The MRQA 2019 Shared Task [29] showed that combining systems like BiDAF, QANet, and BERT improved generalization across out-of-domain datasets.
@@ III-B1 Passage Retrieval Task Dataset
By integrating these diverse resources, we built a more and more representative dataset. The addition enriched the ability of the model to generalize between common and specialized Quranic content, to aid better retrieval performance of verses.
@@ III-D1 Task A: Passage Retrieval Architecture
The original dataset contained 251 questions. To address its limited size, additional questions were sourced from external resources, including Tafseer literature and thematic collections such as plant-related questions in the Qur’an. Each question was paraphrased into two additional variants to ensure linguistic diversity. This process resulted in a dataset comprising over 2,000 unique question formulations. A cleaning phase followed, where duplicate entries were removed and question phrasing was standardized. These steps are illustrated in the data processing module of the architecture (see Figure 1).
@@
Several pre-trained Arabic language models were utilized for fine-tuning. These included AraBERTv02-ARCD, AraELECTRA, CamelBERT- tydi-tafseer, and AraBERTv02-tydi-tafseer. Fine-tuning was carried out with the extended dataset. The training process improved performance in relation to evaluation measures like MAP@10 and MRR@10. Various hyperparameters were explored, including different learning rates, batch sizes, and input sequence lengths. This helped to avoid overfitting and increase generalizability.
@@ III-E Few-Shot Strategy and Prompt Engineering
To optimize the performance of API-based language models like Gemini and DeepSeek, we employed a few-shot prompting approach. This involves adding carefully selected examples to the prompt, helping the model understand the task.
@@ IV-D Discussion
This paper discussed the principal issues in Quranic Question Able to answer passage and span extraction questions. Increasing the dataset from 251 to over 665 distinct questions, alongside paraphrasing, were essential to improving model generality and retrieval accuracy. Although fine-tuning the models for mrc task had limited success because due to data paucity, shifting to instruction-tuned API models like Gemini and DeepSeek. These performed better by few-shot prompt without further training. Ensembling these models continued to improve stability and performance throughout both responsibilities. Nevertheless, the use of APIs presents reproducibility and transparency concerns. Future work should prioritize fully open, reproducible systems, leveraging Graph-based passage linking and retrieval-augmented generation tion for facilitating deeper semantic reasoning across linked Quranic verses
@@ V Conclusion
Future studies would focus on enhancing access to Islamic Understanding. In our upcoming research tasks, we intend to utilize Graph Neural GNNs for representing deeper semantic relationships between passages, enhancing retrieval precision and context-aware answer extraction.

### v2

@@ I Introduction
The Holy Qur’an, revealed over 1,400 years ago, remains the primary source of guidance for over 1.8 billion Muslims worldwide. Beyond its religious significance, the Qur’an represents a masterpiece of Classical Arabic literature, containing profound linguistic, historical, and ethical insights that continue to be studied by scholars across multiple disciplines [1]. However, accessing and understanding specific Quranic content in response to natural language queries presents significant computational challenges that have only recently begun to be addressed by the natural language processing community [2, 3].
Question Answering (QA) systems have emerged as a cornerstone technology for information access across diverse domains, from general knowledge retrieval to specialized applications in legal, medical, and educational contexts [4, 5]. The development of effective QA systems for religious texts, particularly the Qur’an, represents both a meaningful application with substantial social impact and a challenging technical problem that pushes the boundaries of current Arabic NLP capabilities [2, 6].
Quranic Question Answering presents unique challenges that distinguish it from conventional QA tasks. First, the linguistic complexity of Classical Arabic—with its rich morphology, complex syntax, and archaic vocabulary—poses significant difficulties for models primarily trained on Modern Standard Arabic or contemporary dialects [7]. Second, the semantic richness and multi-layered meanings inherent in Quranic verses require sophisticated understanding that goes beyond surface-level text matching. Third, the limited availability of large-scale, high-quality annotated datasets for Quranic content constrains the development and evaluation of machine learning models [8, 9, 10]. Finally, the cultural and religious sensitivity of the domain requires both precision and humility. Quranic interpretation has traditionally been reserved for scholars, and AI outputs risk being mistaken for authoritative tafsīr, potentially misleading users or fueling sectarian disputes. To avoid this, systems should present answers as informational aids rather than interpretive rulings [11].
Recent advances in transformer-based language models have shown promising results for Arabic NLP tasks [14, 15, 12, 13]. However, these models face particular challenges when applied to Classical Arabic texts, where the linguistic gap between training data and target domain significantly impacts performance [16]. While ensemble methods have proven effective in improving robustness across various NLP tasks [17], their application to low-resource, high-complexity domains like Quranic QA remains underexplored [18]. Similarly, the emergence of instruction-tuned large language models offers new possibilities for few-shot learning in specialized domains, yet their effectiveness for Classical Arabic text understanding has not been thoroughly investigated [19, 20, 21].
@@
This paper proposes a two-stage framework combining dataset expansion with tailored modeling: ensemble methods improve retrieval, while instruction-tuned LLMs with prompting enhance extraction.
@@ II-C Ensemble Approaches
Ensemble methods improve QA systems, as shown in MRQA 2019 [17] where combining BiDAF, QANet,and BERT enhanced eneralization. However, Quranic QA presents unique challenges since retrieving isolated verses (āyāt) risks unintended doctrinal weighting, as meaning emerges from interconnected passages within tafsīr traditions [31]. Our system retrieves thematic passages rather than individual verses, preserving coherence while maintaining ensemble benefits.
@@ III-B1 Passage Retrieval Task Dataset
By integrating these diverse resources, we built a more and more representative dataset. The addition enriched the ability of the model to generalize between common and specialized Quranic content, to aid better retrieval performance of [36, 37].
@@ III-D1 Task A: Passage Retrieval Architecture
The original 251-question dataset was expanded using external resources including Tafseer literature and thematic collections. Each question was paraphrased into two variants for linguistic diversity, yielding over 2,000 unique formulations. The dataset was then cleaned by removing duplicates and standardizing phrasing (Figure 2).
@@
Four pre-trained Arabic models were fine-tuned: AraBERTv02-ARCD, AraELECTRA, CamelBERT-tydi-tafseer, and AraBERTv02-tydi-tafseer. Training on the extended dataset improved MAP@10 and MRR@10 performance. Hyperparameter optimization explored learning rates, batch sizes, and sequence lengths to prevent overfitting and enhance generalizability.
@@ III-E Few-Shot Strategy and Prompt Engineering
To optimize the performance of API-based language models like Gemini and DeepSeek, we employed a few-shot prompting approach. This involves adding carefully selected examples to the prompt, helping the model understand the task [40].
@@ IV-D Discussion
This paper discussed the principal issues in Quranic Question Able to answer passage and span extraction questions. Increasing the dataset from 251 to over 665 distinct questions, alongside paraphrasing, were essential to improving model generality and retrieval accuracy. Although fine-tuning the models for mrc task had limited success because due to data paucity, shifting to instruction-tuned API models like Gemini and DeepSeek. These performed better by few-shot prompt without further training. Ensembling these models continued to improve stability and performance throughout both responsibilities. Nevertheless, the use of APIs presents reproducibility and transparency concerns. Future work should focus on open, reproducible systems using graph-based linking and retrieval-augmented generation for deeper semantic reasoning across Quranic verses.
@@ V Conclusion
Future work will enhance access to Islamic understanding by employing Graph Neural Networks (GNNs) to model deeper semantic links between passages, improving retrieval accuracy and context-aware answer extraction.
Acknowledgment
This publication was supported by Qatar University High Impact Grant, Number LAW: 775. The findings achieved herein are solely the responsibility of the authors.